### v2

@@ Abstract.
Despite the remarkable capabilities of Large Language Models (LLMs) in various NLP tasks, they remain vulnerable to hallucinations due to their limited parametric knowledge and lack of domain-specific expertise. Retrieval-Augmented Generation (RAG) addresses this challenge by incorporating external document retrieval to augment the knowledge base of LLMs. In this approach, RAG retrieves document chunks from an external corpus in response to a query, which are then used as context for the downstream language model to generate an answer. However, these retrieved knowledge sources often include irrelevant or erroneous information, undermining the effectiveness of RAG in downstream tasks. To overcome this limitation, we introduce a compact, efficient, and plugable module designed to refine external knowledge sources before feeding them to the generator. The module reconstructs retrieved content by extracting the most relevant and supportive information and reorganizing it into a concise, query-specific format. Through a three-stage training paradigm—comprising supervised fine-tuning, contrastive multi-task learning, and reinforcement learning-based alignment—it prioritizes critical knowledge and aligns it with the generator’s preferences. This method enables LLMs to produce outputs that are more accurate, reliable, and contextually appropriate.
@@ 1. Introduction
Large language models (LLMs) have demonstrated extraordinary versatility in various natural language processing (NLP) tasks, subsuming pipelines that were originally tailormade for each task. Despite being trained on massive text corpora, LLMs still face memory-related challenges such as out-of-date and out-of-domain knowledge, and they occasionally hallucinate non-factual or non-sensical content (Zhou et al., 2021; Maynez et al., 2020). To enhance the accuracy and reliability of LLM-generated outputs, retrieval-augmented generation (RAG) has emerged as a promising solution for knowledge-intensive tasks (Zhu et al., 2023; Gao et al., 2023b; Li et al., 2022) (e.g open-domain question and answering). RAG systems typically follow a “retrieve-then-generate” paradigm (Shao et al., 2023), where a retriever identifies relevant information from an external corpus and uses this information to augment context in constituting the input to a generative model (i.e, generator), thus yielding an improved answer.
Figure 1. An example illustrating vanilla RAG versus RAG with Oreo highlights the impact of redundant and scattered information within the retrieved document chunks. In the vanilla RAG setup, even the retrieved chunks contain contextually relevant information and the correct answer, but the presence of distractions misguides the downstream LM to generate incorrect answers. In contrast, Oreo effectively captures the essential evidence and reconstructs the context, enabling the LM to generate accurate and correct responses.
@@
Despite its promise, a vanilla RAG system usually comes with shortcomings that can hinder its its effectiveness. One major issue is semantic dissonance between the user query, the retriever, and the generator. This occurs when the retrieved documents, while semantically or contextually related to the topic, fail to directly address the query, leading to suboptimal answers (Cuconasu et al., 2024; Wu et al., 2024a). Another challenge pertains to the presence of noise, i.e., misleading, redundant, distracting, or even erroneous information within the retrieved documents. Such noise can misguide the generator, resulting in inaccurate or incoherent answers (Sun et al., 2023a; Shi et al., 2023). For complex tasks that necessitate reasoning across multiple documents, the generator often struggles to correlate dependencies and relationships between them (BehnamGhader et al., 2022), leading to reasoning errors. For example, as illustrated in Figure 1, the correct answer is derived from two different document chunks, but the generator in the vanilla RAG pipeline fails to capture their connections due to the presence of redundant and distracting information. Additionally, RAG systems are prone to the “lost-in-the-middle”(Liu et al., 2023) dilemma, where LMs exhibit a tendency to prioritize information presented at the beginning and end of an input sequence, while paying less attention to the middle. Finally, the lack of joint optimization between the retriever and the generator exacerbates issues such as knowledge inconsistencies (Wang et al., 2023b) or knowledge conflicts (Xu et al., 2024a), which prevent the generator from producing accurate and contextually appropriate responses as the retrieved knowledge fails to adequately support the generation.
@@
To address these challenges, many solutions have been proposed in prior research. Techniques such as query decomposition (Chan et al., 2024; Kim et al., 2023), query rewriting (Wang et al., 2023c; Tan et al., 2024; Chan et al., 2024; Ma et al., 2023b), and query expansion (Lei et al., 2024) aim to improve retriever performance by refining or enriching the input queries. Some studies have integrated rerankers (Yoran et al., 2023; Yu et al., 2023; Nogueira et al., 2019) into retrieval systems, which reorder and prioritize the most relevant documents to ensure that the most pertinent information is provided to the generator. These works attempt to optimize the context on the passage level and largely ensure relevance with the query, but they still face challenges in maintaining comprehensive attention to the nuanced, finer-grained details of query-specific information. Further advancements have been made in noise and redundancy exclusion. For example, filters based on lexical and information-theoretic approaches have been developed to identify and preserve useful content while directly eliminating less relevant information (Wang et al., 2023a; Li et al., 2023; Jiang et al., 2024). Summarization techniques (Xu et al., 2024b) have been developed to synthesize and condense query-focused information from retrieved documents, leveraging extraction or abstraction methods. Compression techniques (Chevalier et al., 2023; Cao et al., 2024a) extend this functionality by generating summary vectors that encode essential information for downstream tasks. While these methods improve efficiency, they do not align the retriever and generator in a manner that guarantees effective collaboration, which often result in knowledge gaps and consequently incorrect or suboptimal generation. From a training perspective, concurrent (Guu et al., 2020; Lin et al., 2023; Zamani and Bendersky, 2024; Izacard et al., 2022) or asynchronous (Zhang et al., 2024b; Shi et al., 2024) training of retrievers and generators is a widely adopted strategy to improve their interaction and collaboration (Guu et al., 2020; Borgeaud et al., 2022). Although such techniques foster synergistic improvements, they can be computationally expensive and often require large amounts of annotated data to achieve optimal results.
@@
Therefore, in this work, we introduce Oreo , a cOntext REcOnstructor designed to enhance the performance of RAG systems on knowledge-intensive tasks by optimizing the quality of context and mitigating knowledge inconsistencies. Oreo is implemented in a plug-and-play manner, functioning as an intermediary module between the retriever and the generator. It receives original document chunks from the retriever and produces refined context tailored for the generator. Instead of merely extracting critical tokens from the chunks, Oreo reorganizes them and generates condensed query-aware summaries. Additionally, Oreo synergizes the reconstructed context with the generator’s behavior of knowledge acquisition, ultimately leading to more accurate and contextually relevant answers. To equip Oreo with robust capabilities for extracting supporting evidence from lengthy contexts and maximizing the accuracy of query responses, we adopt a three-stage fine-tuning paradigm: Stage 1 - supervised fine-tuning (SFT). Oreo is trained to acquire information extraction capabilities by leveraging annotated datasets and knowledge distilled from LLMs. Stage 2 - contrastive learning multitask learning (CML). Oreo enhances its generalizability and identification of self-errors during context reconstruction. Stage 3 - reinforcement learning (RL). The outputs of Oreo are further optimized and aligned with the knowledge demand of downstream generator. In this stage, Oreo learns by leveraging generator’s feedback. To minimize training costs, Oreo utilizes a relatively lightweight language model as its backbone, and keeps the retriever and generator frozen as black-box components. This design ensures modularity and efficiency.
@@
    We propose enhancing the RAG by introducing a ”retrieve-reconstruct-then-generate” paradigm, offering a novel perspective on refining retrieved content for improved integration of external knowledge in RAG. Oreo overcomes the lack of contextual integration among fragmented chunks in vanilla RAG by extracting subtle relations from scattered facts, and transforming redundant context into a concise context.
@@
    We demonstrate Oreo ’s improved performance and reduced token length for both single-hop QA tasks (PopQA (Mallen et al., 2023), NaturalQuestion (NQ) (Kwiatkowski et al., 2019), TriviaQA (TQA) (Joshi et al., 2017), and multi-hop QA tasks (HopotQA (Yang et al., 2018), 2WikiMultiHopQA (Ho et al., 2020)). Overall, Oreo improves an average of 6.87% performance while reducing the input token length for generator by 12.87x.
@@ 2. Methodology
We begin with a quick primer on general retrieval-augmented generation (RAG) and formulate our problem (§2.1), followed by an overview of the proposed method (§2.2) and details of each step (§2.3, §2.4, §2.5 and §2.6).
2.1. Problem Formulation
A typical RAG system comprises of two primary components that work in tandem: the retriever $\mathcal{R}$ identifies and retrieves top-k document chunks $\mathcal{D}=\{d_{\_}{1},d_{\_}{2},...,d_{\_}{k}\}$ from an external knowledge base based on their relevance to a given query q; the generator $\mathcal{G}$ then produces the final answer for q by conditioning on the combination of $\mathcal{D}$ and query q, formally expressed as $y=\mathcal{G}(\mathcal{D},q)$. However, the performance of such a general pipeline is compromised by reasons we indicated in §1. Therefore, we propose Oreo to reconstruct context by extracting the most supportive evidence from $\mathcal{D}$, and producing a concise, query-aware context $\mathcal{C}$ that aligns with the knowledge acquisition mechanics and preference of $\mathcal{G}$. An ideal $\mathcal{C}$ should be produced after Oreo thoroughly understands $\mathcal{D}$, identifies essential entities and facts, establishes their relations, and retains only the necessary information for $\mathcal{G}$ to effectively answer q. This process goes beyond plain information extraction, as it involves organizing and synthesizing content into a coherent and query-specific context. Therefore, we formulate the context reconstruction task as a text generation problem.
@@ 2.2. Method Overview
Our method extends the standard RAG paradigm from “retrieve-then-generate” to “retrieve-reconstruct-then-generate”. Specifically, we train a text generation model _$\mathcal{M}_{\_}{\theta}$, parameterized by $\theta$ to map the retrieved documents $\mathcal{D}$ into a refined context c that enables the downstream generator $\mathcal{G}$ to produce more accurate answers for an input query: $c=f_{\_}{\mathcal{M}_{\_}{\theta}}(\mathcal{D},q)$. The training of _$\mathcal{M}_{\_}{\theta}$ involves three stages: 1. _$\mathcal{M}_{\_}{\theta}$ is trained to learn the transformation from original documents to refined context using annotated datasets (§2.4). 2. Self-generated samples are incorporated to enhance the model’s ability to recognize and correct its own errors, thereby improving robustness and generalization (§2.5). 3. The reconstructed context is aligned with the generator’s knowledge acquisition process by incorporating feedback from $\mathcal{G}$2.6). However, obtaining an annotated dataset with refined context for SFT is challenging. Drawing inspiration from prior work (Balachandran et al., 2022), we replace human annotation with advanced LLMs to generate high-quality synthetic oracle training data (§2.3). An overview of the framework is depicted in Figure 2.
@@ 2.3. Data Collection and Curation
Data collection. To train Oreo during the SFT stage, an annotated dataset containing context with the most supportive evidence from retrieved document chunks is crucial. Such context should be query-specific, answer-aware, grounded in retrieved chunks, and structured as a rationale chain capable of deriving the correct answer. However, such datasets are not readily available, and manually annotating evidence for each query is time-consuming and labor-intensive. Fortunately, contemporary LLMs have exhibited impressive instruction learning capabilities to extract useful information (Dagdelen et al., 2024) and generate high-quality reasoning steps (Wei et al., 2022) even in few-shot settings (Brown et al., 2020). In this work, we elicit such capability from more advanced LLMs to our relatively smaller model Oreo , through generating a high-quality reasoning dataset using LLMs and utilizing it as ”gold context” to train Oreo . Specifically, given a query and corresponding retrieved document chunks, we first prompt Llama-3-8B-Instruct (Touvron et al., 2023) to extract key entities and events from $\mathcal{D}$, and generate detailed rationales to answer the query. Since we prioritize the information extraction capability of Oreo during the SFT stage, to ensure reliability and minimize hallucinations, we construct the gold training dataset solely from query-document pairs where the ground-truth answer is explicitly present within the retrieved chunks.
Bootstrapping. For queries where the generated reasoning fails to include the ground truth (despite it being present in the retrieved chunks), we bootstrap Llama3 by providing the correct answer and iteratively reprompting it to perform generation. Such an iterative process allows Llama3 to reason backwards and learn to generate rationale chains that support the correct answer. This bootstrapping is inspired by (Zelikman et al., 2022; Wei et al., 2024). The prompts and demonstrations used for gold context generation and boostrapping are deatiled in Appendix.
Data curation. Accurate extraction of supporting evidence and reasoning from query to answer is essential for training _$\mathcal{M_{\_}{\theta}}$. To eliminate hallucination and ensure the quality of learning, we conduct data curation by applying the following rules. 1. Ground Truth Alignment. We retain query-context pairs where the generated context from Llama-3 contains ground truth answers. 2. Entity and Event Consistency. We extract sets of entities and events from both the original documents and the Llama3-generated context. Instances are kept only if the entities and events extracted from the generated context (_$\mathcal{E}_{\_}{gen}$) are a subset of those present in the original documents (_$\mathcal{E}_{\_}{ori}$). By following these steps, the refined context generated by Llama-3 is treated as ”gold context” for training _$\mathcal{M_{\_}{\theta}}$.
@@ 2.6. Reinforcement Learning Alignment
Reward estimation. The downstream generator $\mathcal{G}$ serves as a reward model, the generation of Oreo by policy $\pi_{\_}{\theta_{\_}{t}}$ is passed to $\mathcal{G}$ with query $q$ to generate the answer $y$. When the end of sentence (e.g ¡EOS¿) token is generated, the corresponding reward $R_{\_}{t}$ is obtained by comparing the generated answer $y$ with ground truth answer $y_{\_}{gold}$, which is measured by the ROUGE score $R_{\_}{t}=ROUGE(y,y_{\_}{gold})$. However, $\mathcal{G}$ generates answers only after completing all tokens, but (3) updates every action step. To address this, we incorporate a token-level weighting mechanism (Yang et al., 2023). Considering that a token $t$ with higher generation probability deemed more critical by the current policy. Consequently, the token’s contribution to the final reward is proportionally adjusted. We estimate the reward at each step $t$ using the formulation:
(4) $R_{\_}{t}=ROUGE(y,y_{\_}{gold})*log\_softmax(e^{\pi_{\_}{\theta}(a_{\_}{t}|s_{\_}{t})})$
Since the rewards estimated by (4) are sequence-level and sparse, following (Wu et al., 2021) we regularize the reward function using a token-level KL penalty to prevent the model from deviating too far from the initialized LM. The final regularized reward estimation is:
@@
Table 1. Dataset statistics, retrievers and evaluation metrics.
Dataset # Train (k) # Test (k) Retriever Precision@5 Task Metric
PopQA 6.5 1.4 Contriver 0.287 Extractive single-hop QA EM
NaturalQuestions 28.3 3.6 DPR 0.33 Extractive single-hop QA EM
TriviaQA 30.1 11.3 Contriver 0.43 Extractive single-hop QA EM
HotpotQA 20.7 5.6 Contriver 0.137 Abstractive multi-hop QA F1
2WikiMultiHopQA 20.7 12.6 BM25 0.07 Abstractive multi-hop QA F1

### v3

@@ Abstract.
Retrieval-Augmented Generation (RAG) aims to augment the capabilities of Large Language Models (LLMs) by retrieving and incorporate external documents or chunks prior to generation. However, even improved retriever relevance can brings erroneous or contextually distracting information, undermining the effectiveness of RAG in downstream tasks. We introduce a compact, efficient, and pluggable module designed to refine retrieved chunks before using them for generation. The module aims to extract and reorganize the most relevant and supportive information into a concise, query-specific format. Through a three-stage training paradigm—comprising supervised fine-tuning, contrastive multi-task learning, and reinforcement learning-based alignment—it prioritizes critical knowledge and aligns it with the generator’s preferences. This approach enables LLMs to produce outputs that are more accurate, reliable, and contextually appropriate.
@@ 1. Introduction
Large language models (LLMs) have demonstrated remarkable versatility across a wide spectrum of natural language processing (NLP) tasks, subsuming pipelines that were originally tailormade for each task. Despite being trained on massive text corpora, LLMs still face memory-related challenges such as out-of-date and out-of-domain knowledge, and they occasionally hallucinate non-factual or non-sensical content (Zhou et al., 2021; Maynez et al., 2020). To enhance the accuracy and reliability of LLM-generated outputs, retrieval-augmented generation (RAG) has emerged as a promising solution for knowledge-intensive tasks (Zhu et al., 2023; Gao et al., 2023b; Li et al., 2022) (eg., open-domain question and answering). RAG systems typically follow a “retrieve-then-generate” paradigm (Shao et al., 2023), where a retriever identifies relevant information from an external corpus and uses this information to augment context in constituting the input to a generative model (ie., generator), thus yielding an improved answer.
Despite its promise, a vanilla RAG system usually comes with shortcomings that can hinder its effectiveness. One major issue is semantic dissonance between the user query, the retriever, and the generator. This occurs when the retrieved documents, while semantically or contextually related to the topic, fail to directly address the query, leading to suboptimal answers (Cuconasu et al., 2024; Wu et al., 2024a). Another challenge pertains to the presence of noise, ie. misleading, redundant, distracting, or even erroneous information within the retrieved documents. Such noise can misguide the generator, resulting in inaccurate or incoherent answers (Sun et al., 2023a; Shi et al., 2023). For complex tasks that necessitate reasoning across multiple documents, the generator often struggles to correlate dependencies and relationships between them (BehnamGhader et al., 2022), leading to reasoning errors. For example, as illustrated in Figure 1, the correct answer while present in the retrieved chunks is not captured by the vanilla RAG process. Additionally, RAG systems are prone to the “lost-in-the-middle”(Liu et al., 2023b) dilemma, where LLMs exhibit a tendency to prioritize information presented at the beginning and end of an input sequence, while paying less attention to the middle. Finally, the lack of joint optimization between the retriever and the generator exacerbates issues such as knowledge inconsistencies (Wang et al., 2023b) or knowledge conflicts (Xu et al., 2024b), which prevent the generator from producing accurate and contextually appropriate responses as the retrieved knowledge fails to adequately support the generation.
Figure 1. An example comparing vanilla RAG versus RAG with Oreo highlights the impact of redundant and scattered information within the retrieved document chunks. In the vanilla RAG setup, even though the retrieved chunks contain contextually relevant information to the query, the presence of distractions and redundancy misleads the downstream LLM, causing it to misinterpret temporal dependencies and generate an incorrect answer. In contrast, Oreo effectively captures the essential evidence and reconstructs the context, leading to accurate and correct responses.
@@
To address these challenges, many solutions have been proposed in prior research. Techniques such as query decomposition (Chan et al., 2024; Kim et al., 2023), query rewriting (Wang et al., 2023c; Tan et al., 2024; Chan et al., 2024; Ma et al., 2023b), and query expansion (Lei et al., 2024) aim to improve retriever performance by refining or enriching the input queries. Some studies have integrated rerankers (Yoran et al., 2023; Yu et al., 2023; Nogueira et al., 2019) into retrieval systems, which reorder and prioritize the most relevant documents to ensure that the most pertinent information is provided to the generator. These works attempt to optimize the context on the passage level and largely ensure relevance with the query, but they still face challenges in maintaining comprehensive attention to the nuanced, finer-grained details of query-specific information.
Further advancements have been made in noise and redundancy exclusion. For example, filters based on lexical and information-theoretic approaches have been developed to identify and preserve useful content while directly eliminating less relevant information (Wang et al., 2023a; Li et al., 2023; Jiang et al., 2024). Summarization techniques (Xu et al., 2024c) have been developed to synthesize and condense query-focused information from retrieved documents, leveraging extraction or abstraction methods. Compression techniques (Chevalier et al., 2023; Cao et al., 2024a; Cheng et al., 2024; Yoon et al., 2024; Jiang et al., 2023; Pan et al., 2024; Jiang et al., 2024) extend this functionality by generating summary vectors that encode essential information for downstream tasks. While these methods improve efficiency, they do not align the retriever and generator in a manner that guarantees effective collaboration, which often result in knowledge gaps and consequently incorrect or suboptimal generation. From a training perspective, concurrent (Guu et al., 2020; Lin et al., 2023; Zamani and Bendersky, 2024; Izacard et al., 2022) or asynchronous (Zhang et al., 2024b; Shi et al., 2024) training of retrievers and generators is a widely adopted strategy to improve their interaction and collaboration (Guu et al., 2020; Borgeaud et al., 2022). Although such techniques foster synergistic improvements, they can be computationally expensive and often require large amounts of annotated data to achieve optimal results.
@@
In this work we introduce Oreo , a cOntext REcOnstructor designed to enhance the performance of RAG systems on knowledge-intensive tasks by optimizing the quality of context and mitigating knowledge inconsistencies. Oreo is implemented in a plug-and-play manner, functioning as an intermediary module between the retriever and the generator. It receives document chunks from the retriever and produces refined context tailored for the generator. Instead of merely extracting critical tokens from the chunks, Oreo reorganizes them and generates condensed query-aware summaries. Additionally, Oreo synergizes the reconstructed context with the generator’s behavior of knowledge acquisition, ultimately leading to more accurate and contextually relevant answers.
@@
    We propose enhancing the RAG by introducing a “retrieve-reconstruct-then-generate” paradigm, offering a novel perspective on refining retrieved content for improved integration of external knowledge in RAG. Oreo overcomes the lack of contextual integration among fragmented chunks in vanilla RAG by extracting subtle relations from scattered facts, and transforming redundant context into a concise context.
@@
    We demonstrate Oreo ’s efficiency, effectiveness and robustness for both single-hop QA tasks (PopQA (Mallen et al., 2023), NaturalQuestion (NQ) (Kwiatkowski et al., 2019), TriviaQA (TQA) (Joshi et al., 2017)), and multi-hop QA tasks (HopotQA (Yang et al., 2018), 2WikiMultiHopQA (Ho et al., 2020)). On average, Oreo contributes 5.115% downstream performance while reducing the input token length for generator by 12.87x.
@@ 2. Methodology
A typical RAG system comprises of two primary components that work in tandem: the retriever $\mathcal{R}$ identifies and retrieves top-k document chunks $\mathcal{D}=\{d_{\_}{1},d_{\_}{2},...,d_{\_}{k}\}$ from an external knowledge base based on their relevance to a given query q; the generator $\mathcal{G}$ then produces the final answer for q by conditioning on the combination of $\mathcal{D}$ and query q, formally expressed as $y=\mathcal{G}(\mathcal{D},q)$. However, the performance of such a general pipeline is compromised by reasons we indicated in §1. Therefore, we propose Oreo to reconstruct context by extracting the most supportive evidence from $\mathcal{D}$, and producing a concise, query-aware context $\mathcal{C}$ that aligns with the knowledge acquisition mechanics and preference of $\mathcal{G}$. An ideal $\mathcal{C}$ should be produced after Oreo identifies essential entities and facts from $\mathcal{D}$, establishes their relations, and retains only the necessary information for $\mathcal{G}$ to effectively answer q. This process goes beyond plain information extraction, as it involves organizing and synthesizing content into a coherent and query-specific context. Therefore, we formulate the context reconstruction task as (itself) a text generation problem.
@@ 2.1. Method Overview
Our method extends the standard RAG paradigm from “retrieve-then-generate” to “retrieve-reconstruct-then-generate”. Specifically, we train a text generation model _$\mathcal{M}_{\_}{\theta}$, parameterized by $\theta$ to map the retrieved documents $\mathcal{D}$ into a refined context c that enables the downstream generator $\mathcal{G}$ to produce more accurate answers for an input query: $c=f_{\_}{\mathcal{M}_{\_}{\theta}}(\mathcal{D},q)$. The training of _$\mathcal{M}_{\_}{\theta}$ involves three stages: 1. _$\mathcal{M}_{\_}{\theta}$ is trained to learn the transformation from original documents to refined context using annotated datasets (§2.3). 2. Self-generated samples are incorporated to enhance the model’s ability to recognize and correct its own errors, thereby improving robustness and generalization (§2.4). 3. The reconstructed context is aligned with the generator’s knowledge acquisition process by incorporating feedback from $\mathcal{G}$2.5). However, obtaining an annotated dataset with refined context for SFT is challenging. Drawing inspiration from prior work (Balachandran et al., 2022), we replace human annotation with advanced LLMs to generate high-quality synthetic oracle training data (§2.2). An overview of the framework is depicted in Figure 2.
@@ 2.2. Data Collection and Curation
Data collection. To train Oreo during the SFT stage, an annotated dataset containing context with the most supportive evidence from retrieved document chunks is crucial. Such context should be query-specific, answer-aware, grounded in retrieved chunks, and structured as a rationale chain capable of deriving the correct answer. However, such datasets are not readily available, and manually annotating evidence for each query is time-consuming and labor-intensive. Fortunately, contemporary LLMs have exhibited impressive instruction learning capabilities to extract useful information (Dagdelen et al., 2024) and generate high-quality reasoning steps (Wei et al., 2022) even in few-shot settings (Brown et al., 2020). In this work, we elicit such capability from more advanced LLMs to our relatively smaller model Oreo , through generating a high-quality reasoning dataset using LLMs and utilizing it as “gold context” to train Oreo . Specifically, given a query and corresponding retrieved document chunks, we first prompt Llama3-8B-Instruct (Touvron et al., 2023) to extract key entities and events from $\mathcal{D}$, and generate detailed rationales to answer the query. Since we prioritize the information extraction capability of Oreo during the SFT stage, to ensure reliability and minimize hallucinations, we construct the gold training dataset solely from query-document pairs where the ground-truth answer is explicitly present within the retrieved chunks.
Bootstrapping. For queries where the generated reasoning fails to include the ground truth (despite it being present in the retrieved chunks), we bootstrap Llama3 by providing the correct answer and iteratively reprompting it to perform generation. Such an iterative process allows Llama3 to reason backwards and learn to generate rationale chains that support the correct answer. This bootstrapping process is inspired by (Zelikman et al., 2022; Wei et al., 2024). The prompts and demonstrations used for gold context generation and boostrapping are provided in Appendix A.
Data curation. Accurate extraction of supporting evidence and reasoning from query to answer is essential for training _$\mathcal{M_{\_}{\theta}}$. To eliminate hallucination and ensure the quality of learning, we conduct data curation by applying the following rules. 1. Ground Truth Alignment. We retain query-context pairs where the generated context from Llama3 contains ground truth answers. 2. Entity and Event Consistency. We extract sets of entities and events from both the original documents and the Llama3-generated context. Instances are retained only if the entities and events extracted from the generated context (_$\mathcal{E}_{\_}{gen}$) are a subset of those present in the original documents (_$\mathcal{E}_{\_}{ori}$). By following these steps, the refined context generated by Llama-3 is treated as “gold context” for training _$\mathcal{M_{\_}{\theta}}$.
@@ 2.5. Reinforcement Learning Alignment
Reward estimation. With the downstream generator $\mathcal{G}$ serving as a reward model, the generation of Oreo by policy $\pi_{\_}{\theta_{\_}{t}}$ is passed to $\mathcal{G}$ with query $q$ to generate the answer $y$. When the end of sentence (eg., ¡EOS¿) token is generated, the corresponding reward $R_{\_}{t}$ is obtained by comparing the generated answer $y$ with ground truth answer $y_{\_}{gold}$, which is measured by the ROUGE score $R_{\_}{t}=ROUGE(y,y_{\_}{gold})$. However, $\mathcal{G}$ generates answers only after completing all tokens, but (3) updates every action step. To address this, we incorporate a token-level weighting mechanism (Yang et al., 2023). Considering that a token $t$ with higher generation probability deemed more critical by the current policy. Consequently, the token’s contribution to the final reward is proportionally adjusted. We estimate the reward at each step $t$ using the formulation:
(4) $R_{\_}{t}=\textrm{ROUGE}(y,y_{\_}{gold})*\log(\textrm{softmax}(e^{\pi_{\_}{\theta})(a_{\_}{t}|s_{\_}{t})})$
Since the rewards estimated by (4) are sequence-level and sparse, following (Wu et al., 2021), we regularize the reward function using a token-level KL penalty to prevent the model from deviating too far from the initialized LM. The final regularized reward estimation is: